OpenStudy (anonymous):

Find the slope of the line perpendicular to -2x-4y=8

OpenStudy (anonymous):

-2x-4y=8 =>2x+4y=-8 => y= -1/2 x - 2 Slope is -1/2

OpenStudy (anonymous):

so that would be M=-1/2